determine the distance from a long straight wire that would produce a Magnetic Field three times that of the Earth's own field upon the application of 60.0 Ampere currents. (assuming that Earth's Magnetic Field is equal to 5x10^(-5)T and \mu _(0)=4\pi x10^(-7)T. m/A).


1
Expert's answer
2022-05-22T14:38:54-0400

"3B=\frac{\mu_0 I}{2\pi r},\implies"

"r=\frac{\mu_0 I}{6\pi B}=8~cm."